;
DateSerial No.Notes
2018-02-19H490Clear BBR H490 desiccant changed and fitted to upper position prior to MACSSIMIZE test flight C083.
2018-02-19H632Red BBR H632 desiccant changed and fitted to upper position prior to MACSSIMIZE test flight C083.
2018-02-19H394Clear BBR H394 desiccant changed and fitted to lower position prior to MACSSIMIZE test flight C083.
2018-02-19H631Red BBR H631 desiccant changed and fitted to lower position prior to MACSSIMIZE test flight C083.
2017-09-22H632H632 upper red BBR removed for cleaning of both the inner and outer dome. Desiccant also changed.
2017-09-22H490H490 upper clear BBR removed for cleaning of both the inner and outer dome. Discovered inner dome had small scratch on the glass so replaced inner dome. New locking screws used. Desiccant also changed.
2017-09-22H631H631 lower red BBR removed for cleaning of both the inner and outer domes. Desiccant also changed.
2017-09-19H394H394 lower clear BBR removed for cleaning of both the inner and outer dome due to white floating bits inside. Desiccant also changed.
2017-07-26H631H631 red BBR returned to the aircraft. No work carried out, just removed to preserve it and now returned for Clarify.
2017-07-26H632H632 red BBR fitted to the aircraft in preparation for Clarify. No work carried out, just removed to preserve it.
2017-06-28H631H631 red BBR removed from the aircraft due to lack of requirement for chemistry config and to preserve it for Clarify.
2017-06-28H632H632 red BBR removed from the aircraft due to lack of requirement for chemistry config and to preserve it for Clarify.
2017-03-10H631H631 fitted prior to start of WINTEX flying at the request of the Met Office. No work done to it, just removed prior to MOYA to preserve it.
2017-03-10H632H632 fitted prior to start of WINTEX flying at the request of the Met Office. No work done to it, just removed prior to MOYA to preserve it.
2017-02-21H632H632 removed from aircraft prior to MOYA to preserve red domes for Clarify.
2017-02-21H631H631 removed from aircraft prior to MOYA to preserve red domes for Clarify.
2017-01-25H392Lower red BBR (H392 with glass dome number RG695) removed from aircraft as developed a fault within BBR. Duncan to investigate.
2017-01-25H631BBR H631 (was the spare) switched to a red dome (RG695) and fitted to aircraft in lower position. All seems fine with numbers in aircraft (see attached).
2016-10-12H392H392 fitted to the aircraft with a new outer dome due to chip on old one.
2016-10-07H394Clear BBR (SN - H394) fitted to lower position.
2016-10-06H631Lower Clear BBR removed as fault discovered on flight B981. Duncan working to solve problem.
2016-10-06H392Lower Red BBR removed as fault discovered on flight B981. Duncan working to solve problem.
2016-09-15H490H490 (upper clear) calibrated and refitted
2016-09-15H632H632 (upper red) calibrated and refitted
2016-09-15H631H631 (lower clear) calibrated and refitted
2016-09-15H392H392 (lower red) calibrated and refitted
2016-09-15H394H394 (spare clear) calibrated
2016-09-12H392Found to be going negative when calibrated on the zero cycle, which was also upsetting the other inputs to the Campbell Scientific CR23X logger, causing a "pedestal" on the other BBRs' zero measurement before going to a "sensible" value. When removed from the calibration rig, the output was found to be going to the neagtive rail (-15V), which was outside the range of the logger. When opened, the circuit board was corroded and a connection wire had come adrift. Board cleaned, wire remade, and other potential dodgy ones also remade.
2016-08-03H631Old desiccant removed and replaced with new desiccant.
2016-08-03H632Old desiccant removed and replaced with new desiccant.
2016-08-03H392Old desiccant removed and replaced with new desiccant.
2016-08-03H490Old desiccant removed and replaced with new desiccant.
2016-08-03H394Old desiccant removed and replaced with new desiccant.
2016-07-03noneFollowing flight B970 UPPER BBR (clear) was removed as condensation has developed within the dome. The dome was removed, cleaned and refitted.
2016-07-03nonePost flight B970 the UPPER BBR (red) was removed, the dome was cleaned and refitted.
2016-06-07H392H392 was fitted replacing H394 which had developed a fault. The calibrations constants that should be applicable from the most recent DECADES calibration file are: X0 X1 CALUP1S 1.159271E+00 2.668560E-02 H392 CALUP2S 3.891574E-01 2.926867E-02 H632 CALLP1S 1.493447E+00 4.304429E-02 H631 CALLP2S 1.547564E+00 5.348979E-02 H490 **ADDITIONAL INFO** added 25/08: The above suggests that H392 (which replaced faulty clear H394) and H632 were in upper positions 1 and 2, and that H631 and H490 were in lower positions 1 and 2 respectively. However – H392 and H632 are both red, and H631 and H490 are both clear – so this doesn’t make much sense (unless the glass was changed in two of them, and this wasn’t recorded). Further info received from AW 20/08: "...We initially did what was entered in the db for simplicity. However in the Indian sun the red instrument with the clear dome on the top over-ranged. So we swapped the top instrument in the clear dome position with the bottom instrument in the red position, also swapping the domes to make the correct config. Jamie and bex were both there and actually did the work..."
2014-05-28noneA fault with the LOWER BBR data logging system was discovered for FLIGHTS B848, B849, B850. The LOWER (upwelling) radiation fluxes for these flights does not exist in a useful form due to the fault. The fault was identified, repaired and tested WORKING for flights B851 onwards.
2014-05-28noneA fault with the LOWER BBR data logging system was discovered for FLIGHTS B848, B849, B850. The LOWER (upwelling) radiation fluxes for these flights does not exist in a useful form due to the fault. The fault was identified, repaired and tested WORKING for flights B851 onwards.
2012-08-02H392Red instrument (H392) fitted with clear dome and put in lower position to provide in-flight testing capability for Lower BBR DLU (DECADES)
2011-07-06noneLower clear BBR cable had been switched with position 3 owing to cable-length issues. The same change had not been made at the top/dlu end, so clear pyranometer data were not being processed correctly. The cables have now been swapped at both ends, Clear BBR data now return to the correct DLU channels.
2011-05-16H631Lower Instrument removed. Dome spotless, cleaned and refitted.
2011-05-16H394Was upper instrument. Dome showing marked dirt accretion on forward face (following transit back from pre-FENNEC). Dome replaced with clean unused one. Alignment of upper cover was displaced clockwise of BBR body by 120 degrees, not known which alignment was used to fit the instrument but this may explain alignment offsets found by Reading from pre-FENNEC data.
2011-05-16H490Lower Red instrument removed unserviceable. Replaced with H490 (normally operated as a clear dome instrument), with red dome fitted.
2010-04-19H394Work done not logged in database
2010-04-19H392Work done not logged in database
2010-04-19H632Work done not entered into database
2010-04-19H631Work done was not previously logged
2010-04-14H631removed aircraft strip and respray.
2010-04-14H394removed aircraft strip and respray.
2010-04-14H392removed aircraft strip and respray.
2010-04-14H632removed aircraft strip and respray.
2009-11-03H392Work done was not entered in database
2009-11-03H632Work done was not entered in database.
2009-09-04H495Not required for TROMPEX.
2009-09-04H632Not required for TROMPEX.
2009-07-28H495Water ingress has damaged circuit boards
2009-07-17H490Dirt, possibly silica gel fragments, visible in inner and outer domes.
2009-07-17H631consts updated in fltcons
2009-04-16H490Cabling error corrected.
2009-03-20H490NULL
2009-03-20H394NULL
2009-03-20H495NULL
2009-03-20H392NULL
2009-03-13H490removed for cal but flight constants not updated until after flight B435
2009-03-13H394removed for cal but flight constants not updated until after flight B435
2009-03-13H631removed for cal but flight constants not updated until after flight B435
2009-03-13H490NULL
2009-03-13H632removed for cal but flight constants not updated until after flight B435
2009-03-13H394NULL
2009-03-13H495NULL
2009-03-13H392NULL
2009-03-13H495removed for cal but flight constants not updated until after flight B435
2009-03-13H392removed for cal but flight constants not updated until after flight B435
2008-10-15H490NULL
2008-10-15H394NULL
2008-10-15H495NULL
2008-10-15H392NULL
2008-10-01H394NULL
2008-10-01H490NULL
2008-10-01H495NULL
2008-10-01H392NULL
2008-08-06H392NULL
2008-08-06H495NULL
2008-06-09H495Not required for OP3
2008-06-09H632Not required for OP3
2007-12-10H490Cable for lower clear and pyrgeometer corrected.
2007-11-27H394NULL
2007-11-27H632NULL
2007-11-12H490NULL
2007-11-12H394NULL
2007-11-12H495NULL
2007-11-12H632NULL
2007-10-12H490Taken to Cardington for intercomparison.
2007-10-11H394Taken to Cardington for intercomparison
2007-10-11H495Domes replaced with clear ones and taken to Cardington for intercomparison.
2007-10-11H632Dome replaced with clear one and taken to Cardington for intercomparison.
2007-06-12H490NULL
2007-06-12H495NULL
2007-05-18H631NULL
2007-05-18H392NULL
2007-04-27H490Ingress of water into outer dome
2007-04-27H631Fitted to replace saturated H490
2007-04-16H490Fitted to replace saturated H631
2007-04-16H631Inner and outer domes full of water, thermopile saturated
2007-04-16H631Both sets of domes removed and thermopile dried using hairdrier on cold
2007-01-17H631NULL
2007-01-17H490Dome was dirty.
2006-11-06H394Clear dome 003
2006-11-06H490Clear dome 001
2006-11-06H632Red dome 002
2006-11-06H392Red dome 001